OpenStudy (anonymous):

Which ordered pair is on the graph of the equation 3x + 7y = 11? A) (1, 2) B) (0,4) C) (6, 1) D) (0, 4)

OpenStudy (anonymous):

A) (1, -2) B) (0, -4) C) (6,-1) D) (0,4) sorry forgot the negatives

OpenStudy (hoblos):

replace the coordinates and try no one is right !!

OpenStudy (campbell_st):

substitute the x value from each ordered pair into the equation to see which y value you get

OpenStudy (anonymous):

Yeah...I tried each one, and none of them work.

OpenStudy (hoblos):

ok now C is right 3(6) + 7(-1) = 18-7 =11

OpenStudy (anonymous):

Hoblos, you have to ADD them...you can't subtract.

OpenStudy (hoblos):

yeah.... 18 + (-7) = 18 - 7 =11

OpenStudy (anonymous):

Ooohh. It's a negative 7?

OpenStudy (hoblos):

when you multiply it by (-1) it becomes negative

OpenStudy (anonymous):

Oh...psh. I knew that. Haha. Thanks for clearin' that up.

OpenStudy (hoblos):

my pleasure XD